# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

### v1

@@ 4 Why Sharp-Eyed Refinement?
In this section, we analyze the limitations of our captioning specialist by proposing two hypotheses and validating them through toy experiments and visualizations. We then introduce our captioning framework to alleviate these issues.
Analysis 1.
@@
(Hypothesis.) We propose a first hypothesis: they fail to focus on essential critical regions (e.g., entities, attributes, relations), instead maintaining a redundant and blurry view of the entire image. This may stem from the current captioning framework, which requires the model to observe the image and generate the entire caption in a single pass. This approach can be likened to asking a human to describe an image in a single glance, without focusing on specific details. (Verification.) To investigate this issue, we analyze which regions the model attends to during the generation of specific words. As detailed in Sec. H.3, the attention maps are visualized by implementing code that references techniques from the API [88]. As illustrated in Fig. 5, the model often exhibits diffuse, global attention and fails to focus on specific regions. These findings underscore the necessity for an approach that guides the captioning process toward more effective attention mechanisms.
Note 2:
@@
Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary for precise description.
Analysis 2.
@@
(Hypothesis.) Eyes Wide Shut [66] shows that visual modality deficiencies-specifically, the CLIP vision encoder [56] yielding similar representations for distinct images (a phenomenon termed visual blindness)-undermine MLLM performance. Building on these findings, we hypothesize that the widely used CLIP features may produce ambiguous representations. (Verification.) We attempt to reconstruct images from their feature embeddings. The quality of the reconstructed images serves as an indicator of how much visual detail the features preserve and, consequently, their effectiveness in supporting captioning tasks. For this experiment, we directly employ the decoder architecture from Masked Autoencoder (MAE) [22]. We keep the CLIP encoder frozen while training only the MAE decoder on the COCO images. Our goal was to visualize how the image is represented within the language model. As illustrated in Fig. 6, the images reconstructed from CLIP embeddings exhibit significant deviations from the original images and appear less clear. This result indicates that such visual representations may not provide sufficiently detailed visual information for captioning.
Note 3:
@@
CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
One More Glance.
Building on the above insights, we propose a novel captioning framework, Sharp-Eyed Refinement. As illustrated in Fig. 1 (b) and (c), our framework is inspired by the human description process. Specifically, the model first generates an initial caption. Then, the same LM, in conjunction with our proposed multimodal connector DeepLens, refines this output. This refinement step mimics a human taking a second glance to revise their initial impression, allowing the model to generate captions with more effective attention mechanisms and fine-grained grounding (i.e., with sharp eyes). The design and functionality of DeepLens are described in the following section.
Our refinement process is similar to ideas explored in LLM research through self-refinement [50; 32; 54], where models improve their outputs by revisiting previous generations. Extending this concept, we apply refinement to captioning models, uniquely integrating visual features for multimodal refinement. To the best of our knowledge, this represents the first attempt of its kind.
@@ 5 Inside Our Framework: DeepLens and Learning Strategy
We aim to mitigate the above limitations by incorporating our proposed DeepLens into the caption refinement process. In particular, unlike prior multimodal connectors [43; 37; 3], DeepLens takes two new inputs: (i) previously generated captions and (ii) multi-layer features from the vision encoder. Below, we examine the significance of this design in detail.
Looks at what matters.
@@
Imagine receiving an image along with the caption “A cat relaxing on a brown chair” and being instructed to refine it. In such a scenario, humans naturally focuses on key elements such as the cat and the chair. We attempt to employ a similar mechanism by providing the previously generated caption as an additional input to guide the refinement process. This design enables the module to identify and attend to the relevant visual regions. Similar to the experiment in Sec. 4, we also inspect the attention maps while the model takes a second glance using the previous captions, as shown in Fig. 5. We observe that the model focuses more precisely on regions corresponding to specific words, indicating that it attempts to look at what matters in the image.
Looks in detail.
@@
To mitigate the coarse-grained visual feature problem, some studies proposed using additional vision towers [65; 66; 27], but this increases computational costs. For example, Interleaved MoF [66] adds 300M parameters for DINOv2 [53]—a substantial 60% increase relative to our 500M parameter model. Instead, we focus on maximizing the utility of the existing vision encoder. We design DeepLens to leverage features extracted from various layers of the vision encoder, as these multi-level features enable more semantically rich encoding of detailed image elements, following FPN [41]. Indeed, the reconstruction experiment conducted in Fig. 6 demonstrates that utilizing these features helps generate clearer images. This implies that the visual embeddings manipulated by DeepLens using multi-level features contain sufficient fine-grained information about the image, supporting the LM in recognizing visual content more accurately.
Note 4:
@@
Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
The architectural design of DeepLens follows a BERT [17] structure that takes two inputs: visual features and projected embeddings of the initial captions. Further details are in Sec. C.
Figure 7: Qualitative results of Our Sharp-Eyed Refinement. More examples are in Sec. E.
Figure 8: To develop our specialist, our model undergoes a 2-step fine-tuning process for initial caption generation and refinement. During the inference stage, the caption is initially generated and then refined, following the green and blue lines, respectively.
Fine-tuning Strategy.
@@
Assume we have a training set $X=\{x_{1},\dots,x_{N}\}$, where each $x_{k}=(i_{k},c_{k})$ consists of an image $i_{k}$ and its corresponding ground-truth caption $c_{k}$; $N$ denotes the total number of such pairs. As illustrated in Fig. 8, we adopt a two-stage fine-tuning strategy. In the first stage, we focus on enabling the model to generate an initial caption by following the training procedure of LLaVA. Specifically, we perform supervised fine-tuning on the MLP connector and the lightweight language model using this dataset.
@@
The second stage aims to facilitate the model’s ability to learn the refinement process effectively. We first consider that directly using the momentary generated captions as initial inputs and the true captions $c_{k}$ as the targets for refinement can be misleading. Suppose our captioning specialist were to initially produce something like ‘a brown table in front of a window’ while the true caption is ‘a cat sitting on a chair’; these two descriptions share little in common and provide no meaningful signal for gradual correction. Training on such disparate pairs would encourage the model to ignore the provided initial caption and generate a wholly new one, rather than learn to refine it. Additional explanations are given in Sec. D.5.
To address this challenge, we construct pseudo-initial captions $\hat{c}_{k}$ that remain close to their corresponding ground-truth $c_{k}$. Specifically, we leverage GPT-4o-mini [1] to possibly introduce small errors into each $c_{k}$, producing slightly altered versions that focus on entities, attributes and relations (e.g., ‘a dog sitting on a chair’). During second-stage fine-tuning, each pair $(i_{k},\hat{c}_{k})$ is fed into DeepLens alongside the language model, and a cross-entropy loss is applied. Our fine-tuning regimen consists of 10 epochs for the first stage and 2 epochs for the second stage. The details and examples of how $\hat{c}_{k}$ is generated are provided in Sec. F.1. The resulting dataset $X$, including the pseudo-initial captions $\hat{c}_{k}$, will be made publicly available.
Rationale for Refinement Fine-tuning.
@@
Let $\pi_{\theta}$ denote the language model. Each pseudo-initial caption $\hat{c}_{k}$ deviates from the ground truth $c_{k}$ at only a few token positions $E_{k}=\{t\mid\hat{c}_{k,t}\neq c_{k,t}\}$. Under the sequence-level cross-entropy objective
@@
gradients are likely to be primarily concentrated on tokens in $E_{k}$, leading to a form of targeted optimization, in which the model retains the correct parts of $\hat{c}_{k}$ while rewriting only the erroneous ones. Writing $\Delta_{k}(\theta)=\log\pi_{\theta}(c_{k}\mid i_{k},\hat{c}_{k})-\log\pi_{\theta}(\hat{c}_{k}\mid i_{k},\hat{c}_{k}),$ we have $\mathcal{L}(\theta)\propto-\mathbb{E}[\Delta_{k}]$; hence, minimizing $\mathcal{L}$ is equivalent to maximizing the expected margin $\Delta_{k}$, thereby directly increasing the likelihood of the refined caption relative to its flawed precursor. In effect, each gradient step encourages DeepLens to focus on visual regions likely responsible for errors in the initial caption and guides the language model to better interpret these refined features, resulting in more accurate captions. We note that this targeted optimization resembles the philosophy of Direct Preference Optimization (DPO) [57], if we consider $c_{k}$ and $\hat{c}_{k}$ as preferred and less-preferred responses. In contrast to DPO, which treats both responses symmetrically during optimization, our method offers a new perspective by assigning them distinct roles as input and target.
Inference Strategy.
Given an image $i_{k}$ and instruction $p$, the model first generates an initial caption $o_{initial}$, storing intermediate ViT [19] features in a buffer $B$ to avoid redundant computation. $o_{initial}$ is then fed into DeepLens and the LM to produce the refined output $o_{refined}$. Although this process can be carried out iteratively, similar to Self-Refine [50] in LLM research, our experiments indicate that it does not provide additional performance gains. We therefore do not adopt a repeated process but anticipate it to be a promising direction for future research, as described in Sec. D.6 and Sec. B.6.
MS COCO [14]
model #params B@4 [55] gain CIDEr [71] gain CLAIR [8] gain GPT [9] gain
@@
ShareGPT4V [11] & DCI [70]
model #params CIDEr [71] gain CAPT [18] gain CLAIR [8] gain GPT [9] gain
@@
GLaMM [60]
model #params CIDEr [71] gain CAPT [18] gain CLAIR [8] gain GPT [9] gain
@@
Table 3: Quantitative evaluation of Sharp-Eyed Refinement (SeR). The results demonstrate the effectiveness of SeR in improving caption quality. We also present an ablation study on the key inputs to DeepLens: ① initial captions and ② multi-layer features.
@@ 6 Empirical Validation
Ablation Study of DeepLens. To assess the contributions of DeepLens, we conduct an ablation study as shown in Tab. 3 and Fig. 7. The results show that our framework yields both quantitative and qualitative improvements in caption quality. The refinement process enhances initial captions by +3.9 and +2.5 CIDEr points for single-sentence and detailed captioning tasks, respectively. While the integration of multi-layer visual features (i.e., with ②) provides modest gains, its effectiveness remains constrained by the limitations of single-pass generation. In fact, our method introduces additional parameters (approximately 50M for DeepLens) and an additional inference step with OPT-125M. However, considering the significant advantages shown in Tab. 9, we believe that the computational overhead introduced by our refinement method is minimal. This measurement was conducted over 100 iterations with a batch size of 1.
@@
Since captioners are expected to serve as foundational components in downstream applications as mentioned in Sec. 2, it is important to verify their utility in practical settings. To this end, we evaluate various captioners on the Long-Range Video Question Answering task introduced in [90]. Following the baseline setup of Zhang et al. [90], each captioner generates captions for multiple frames from videos spanning several minutes to hours. These captions are then fed into an LLM to answer video-related questions. As shown in Tab. 4, our specialist achieves competitive accuracy with significantly fewer parameters and faster inference than large-scale generalists, demonstrating its potential for real-world applications discussed in Fig. 1. The reported time includes both the captioning and LLM inference steps, assuming a 10-minute video processed at the frame level. The small specialists [58; 25], including ours, use weights fine-tuned on the ShareGPT [11] and DCI [70] datasets.
LLM Captioner LLaVA-1.0 [43] BLIP-2 [37] LLavA-1.5 [42] SmallCap* [58] Tag2text* [25] Our specialist + SeR
Qwen2.5 14B [84] #params 7.3B 7.4B 7.3B 450M 900M 450M 500M
@@
Table 4: Evaluation on Long-Range VideoQA. We follow the baseline setup of LVQA [90], replacing the captioner with our specialist and large multimodal generalists. Captions from all frames are aggregated and fed into Qwen2.5-14B to perform the final video question answering.
Note 5:
Despite its significance, the use of lightweight captioners has been underexplored. Our proposed baseline and framework demonstrate the potential of lightweight captioning for on-device applications.

### v2

@@ 4 Why Sharp-Eyed Refinement?
In this section, we conduct a deeper investigation into our captioning specialist, formulating two hypotheses and validating them through toy experiments and visualizations. We then present our captioning framework to address these issues.
@@
Analysis 1.(Hypothesis.) We propose a first hypothesis: they fail to focus on critical regions (e.g., entities, attributes, relations), instead maintaining a blurry view of the entire image. One possible cause lies in the current captioning framework, which forces the model to process the entire image and produce a caption in a single pass. This is analogous to asking a human to describe an image in one glance, without attending to specific details. (Setup.) To investigate this issue, we analyze where in the image the model attends during the generation of specific words. The attention maps are visualized using code adapted from API yu2024attention , with further implementation details provided in Section I.3. (Result.) As illustrated in Figure 5, the model often exhibits diffuse attention and fails to focus on specific regions. These observations motivate the development of approaches that guide the captioning process toward more effective attention mechanisms.
@@
Note 2: Single-pass captioning can lead models to adopt broad attention, neglecting critical visual regions necessary for precise description.
@@
Analysis 2.(Hypothesis.) Eyes Wide Shut eyeswideshut demonstrates that MLLMs suffer from visual deficiencies, as vision encoders such as CLIP CLIP provide ambiguous representations that weaken performance. Based on this, we hypothesize that the widely used CLIP features may not provide sufficiently informative representations for generating accurate captions. (Setup.) Unlike prior studies, we examine this issue by reconstructing images from their visual features. The quality of the reconstructed images indicates how much visual detail the features preserve and, consequently, their effectiveness in supporting captioning tasks. For this experiment, we directly employ the decoder architecture from Masked Autoencoder (MAE) mae . We keep the CLIP encoder frozen while training only the MAE decoder on the MS-COCO images. (Result.) As illustrated in Figure 6, the images reconstructed from CLIP embeddings exhibit significant deviations from the original images and appear less clear. This result indicates that such visual representations may not provide sufficiently detailed visual information for captioning.
@@
Note 3: CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
Our proposed solution.Building on the above insights, we propose a novel captioning framework, Sharp-Eyed Refinement (SeR), illustrated in Figure 1 and Figure 7. The framework is inspired by the human description process, where one first takes an initial glance and then sharpens the eyes to capture finer details. Concretely, the model generates an initial caption that is subsequently refined by the LM through our DeepLens. This refinement process resembles approaches in LLM research such as self-refinement selfrefine ; le2022coderl ; pan2023automatically , where models improve their outputs by revisiting earlier generations. By extending this idea to the multimodal domain, we uniquely integrate visual features into the refinement step, and, to the best of our knowledge, this constitutes the first work in this direction. Further implementation details are provided in the following section.
@@ 5 Inside Our Framework: DeepLens and Learning Strategy
Figure 7: Inference and training pipeline for SeR. The model is trained in two fine-tuning stages: (1) initial caption generation and (2) caption refinement. At inference, both an initial caption and a refined caption are produced, as shown by the solid and dashed lines in the rightmost figure.
In this section, we take a closer look at our framework and its effectiveness. At the core lies DeepLens, a multimodal connector that, unlike prior designs llava ; li2023blip2 ; bai2023qwenvl , leverages two new inputs: (i) previously generated captions and (ii) multi-layer features from the vision encoder. These inputs guide our model to attend to salient regions, enrich visual representations, and enable effective caption refinement.
@@
Look at what matters.Imagine receiving an image along with the caption “A cat relaxing on a brown chair” and being instructed to refine it. In such a scenario, humans naturally focus on key elements like the cat and the chair. We attempt to employ a similar mechanism by providing the initially generated caption as an input for DeepLens. This design enables the module to identify and attend to the relevant visual regions. Similar to the experiment in Section 4, we also inspect the attention maps while the model takes a second glance using the previous captions, as shown in Figure 5. We observe that the model focuses more precisely on regions corresponding to specific words, indicating that it attempts to look at what matters in the image.
@@
Look in detail.To mitigate the limitation of coarse-grained visual features, some studies cambrian ; eyeswideshut ; fromcliptodino proposed using additional vision encoders, but this comes at the cost of additional model parameters. For example, Interleaved MoF eyeswideshut adds 300M parameters for DINOv2 oquab2024dinov —a substantial 60% increase relative to our 500M-parameter model. Rather than introducing additional encoders, we focus on maximizing the utility of the existing vision encoder. To this end, we design DeepLens to leverage multi-layer features, expecting them to provide semantically richer representations that enable more accurate captioning, inspired by FPN fpn . As demonstrated in the reconstruction experiment (Figure 6), utilizing these features yields clearer images. This suggests that the multi-level visual features indeed contain fine-grained information, thereby helping the LM recognize visual content in greater detail.
@@
Note 4: Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
Figure 8: Examples of captions during SeR. Initial outputs occasionally include inaccuracies, which are corrected through the refinement process. More results can be found in Section F.
@@
Fine-tuning strategy.Assume we have a training set $X=\{x_{1},\dots,x_{N}\}$, where each $x_{k}=(i_{k},c_{k})$ consists of an image $i_{k}$ and its corresponding ground-truth caption $c_{k}$; $N$ denotes the total number of such pairs. As illustrated in Figure 7, we adopt a two-stage fine-tuning strategy. In the first stage, we focus on enabling the model to generate an initial caption by following the training procedure of LLaVA. Specifically, we perform supervised fine-tuning on the MLP connector and the lightweight language model using this dataset.
@@
In the second stage, we focus on teaching the model the refinement process. A naïve approach would be to directly use the momentary generated captions as initial inputs and the true captions $c_{k}$ as targets for refinement, but this can be misleading. Suppose our captioning specialist were to initially produce something like ‘a brown table in front of a window’ while the true caption is ‘a cat sitting on a chair’; these two descriptions share little in common and provide no meaningful signal for progressive refinement. We believe that training on such disparate pairs would encourage the model to ignore the provided initial caption and generate a wholly new one, rather than learn to refine it. We provide additional explanations in Section E.5. To address this challenge, we construct pseudo‐initial captions $\hat{c}_{k}$ that remain close to their corresponding ground-truth $c_{k}$. In practice, we leverage GPT-4o-mini achiam2023gpt to slightly alter entities, attributes, or relations within $c_{k}$. For example, given the true caption ‘a cat sitting on a chair’, the pseudo-initial caption might become ‘a dog sitting on a chair,’ preserving the overall structure while introducing minor variations.
Leveraging these pseudo-initial captions, each pair $(i_{k},\hat{c}_{k})$ is fed into DeepLens together with the language model, and optimization is performed using a next-token prediction loss (i.e., supervised fine-tuning). Our fine-tuning schedule consists of 10 epochs for the first stage and 2 epochs for the second stage. Further details of the training setup and hyperparameters are provided in Section I.1, and we describe the generation strategy for $\hat{c}_{k}$ in Section G.1, along with an ablation study on refinement data in Section E.5.
@@
Rationale for refinement fine‑tuning.Let $\pi_{\theta}$ denote the language model. Each pseudo‑initial caption $\hat{c}_{k}$ deviates from the ground truth $c_{k}$ at only a few token positions $E_{k}=\{t\mid\hat{c}_{k,t}\neq c_{k,t}\}$. Under the sequence‑level cross‑entropy objective
@@
gradients are likely to be primarily concentrated on tokens in $E_{k}$, leading to a form of targeted optimization, in which the model retains the correct parts of $\hat{c}_{k}$ while rewriting only the erroneous ones. Writing $\Delta_{k}(\theta)=\log\pi_{\theta}(c_{k}\mid i_{k},\hat{c}_{k})-\log\pi_{\theta}(\hat{c}_{k}\mid i_{k},\hat{c}_{k}),$ we have $\mathcal{L}(\theta)\propto-\mathbb{E}[\Delta_{k}]$; hence, minimizing $\mathcal{L}$ is equivalent to maximizing the expected margin $\Delta_{k}$, thereby directly increasing the likelihood of the refined caption relative to its flawed precursor. In effect, each gradient step encourages DeepLens to focus on visual regions likely responsible for errors in the initial caption and guides the language model to better interpret these refined features, resulting in more accurate captions. We note that this targeted optimization resembles the philosophy of Direct Preference Optimization (DPO) dpo , if we consider $c_{k}$ and $\hat{c}_{k}$ as preferred and less-preferred responses. In contrast to DPO, which treats both responses symmetrically during optimization, our method offers a new perspective by assigning them distinct roles as input and target.
Inference strategy.Given an image $i_{k}$ and instruction $p$, the model generates an initial caption $o_{initial}$, storing intermediate ViT vit features in a buffer $B$ to avoid redundant computation. $o_{initial}$ is then fed into DeepLens and the LM to produce the refined output $o_{refined}$. Unless otherwise specified, we employ a single-step refinement. A discussion and experimental results for a repeated process are presented in Section E.6 and Section C.6, respectively.
MS COCO mscoco
model #params B@4 papineni2002bleu gain CIDEr vedantam2015cider gain CLAIR clair gain GPT mllmjudge gain
@@
ShareGPT4V chen2024sharegpt4v & DCI dci
model #params CIDEr vedantam2015cider gain CAPT dong2024benchmarking gain CLAIR clair gain GPT mllmjudge gain
@@
GLaMM rasheed2024glamm
model #params CIDEr vedantam2015cider gain CAPT dong2024benchmarking gain CLAIR clair gain GPT mllmjudge gain
@@
Table 3: Quantitative results of Sharp-Eyed Refinement (SeR). The results demonstrate the effectiveness of SeR in improving caption quality. We also present an ablation study on the key inputs to DeepLens: ① initial captions and ② multi-layer features.
@@ 6 Empirical Validation
Ablation Study of SeR. The results in Table 3 and Figure 8 demonstrate the effectiveness of SeR in improving caption quality. We further conduct an ablation study on the key inputs to DeepLens—① initial captions and ② multi-layer visual features—to examine their individual contributions. The refinement process improves initial captions by +3.9 and +2.5 CIDEr points for single-sentence and detailed captioning tasks, respectively. Although integrating only multi-layer features (i.e., using ②) yields modest gains, we observe that its effectiveness may be constrained by the single-pass generation mechanism. Moreover, our method introduces approximately 50M additional parameters for DeepLens and one extra inference step with OPT-125M. Nevertheless, as shown in Table 9, the computational overhead remains minimal relative to the substantial improvements achieved. We further validate this efficiency on a resource-constrained device, Jetson Nano, with results reported in Section C.3.
@@
Since captioning models serve as foundational components in downstream applications as mentioned in Section 2, it is important to verify their utility in a practical task. To this end, we evaluate our model on the Long-Range VQA task introduced in zhang2023simple . Following the baseline setup of Zhang et al. zhang2023simple , each captioner generates captions for multiple frames from videos spanning several minutes to hours. These captions are then fed into an LLM to answer video-related questions. As shown in Table 4, our model achieves competitive accuracy with significantly fewer parameters and faster inference than MLLM generalists, demonstrating its potential for real-world applications. The reported time includes both the captioning and LLM inference steps, assuming a 10-minute video processed at the frame level. The small specialists ramos2023smallcap ; huang2023tag2text , including ours, use weights fine-tuned on ShareGPT chen2024sharegpt4v  & DCI dci .
LLM Captioner LLaVA-1.0 llava BLIP-2 li2023blip2 LLavA-1.5 llava15 SmallCap* ramos2023smallcap Tag2text* huang2023tag2text Our specialist + SeR
Qwen2.5 14B yang2024qwen2 #params 7.3B 7.4B 7.3B 450M 900M 450M 500M
@@
Table 4: Evaluation on Long-Range VideoQA. We follow the baseline setup of LVQA zhang2023simple , replacing the captioner with our specialist or MLLM generalists. Captions from all frames are aggregated and fed into Qwen2.5-14B to perform the final video question answering. Higher-quality captions from our specialist contribute to competitive accuracy with significantly fewer parameters.
Note 5: Although lightweight captioners play an important role, they have received limited attention. Our proposed baseline and framework highlight their promise as efficient solutions for on-device applications.